legacy manufacturing has a 10% cost of capital, and has generated the following information on two mutually exclusive projects

Year Project A Project B
0 -$110,000 -$90,000
1 $50,000 $40,000
2
$50,000
$40,000
3
$50,000
$40,000
4
$50,000
$40,000

A compute the net present value (NPV) for project A?

B. Compute the (NPV) for project B?

C. Which project should Legacy adopt? Why?
Answer #1

Project A)

Using NPV formula,

NPV = CF0+CF1/(1+r1)+ CF1/(1+r2)+ ....................+ CF1/(1+rn)

NPV = -110000 + 50000/(1+0.1)1+ 50000/(1+0.1)2+50000/(1+0.1)3+50000/(1+0.1)4

= -110000+50000(1/(1+0.1)1+ 1/(1+0.1)2+1/(1+0.1)3+1/(1+0.1)4)

= -110000+50000(0.909+0.8264+0.751+0.683) = 48470 approx.

Project B)

NPV = -90000 + 40000/(1+0.1)1+ 40000/(1+0.1)2+40000/(1+0.1)3+40000/(1+0.1)4

= 36776 approx.

C) Since the NPV of the project A is more, we should adopt project A